# Advancing Image Super-resolution Techniques in Remote Sensing: A Comprehensive Survey

### v2

@@ 3.1.3 Regression-based
The regression-based RSISR methods learn the function mappings between LR and HR images, which regards super-resolution as a regression problem. More specifically, these methods usually extract features from LR images based on paired datasets to train regression models, and then use the trained models to predict and reconstruct HR images. Typically, regression-based methods consist of three main steps: regression model training, projecting of LR images to HR image space, and final super-resolution result generation, as shown in Fig. 5 (bottom row).
@@
In the past two decades, various regression-based super-resolution methods have been proposed, including kernel regression-based methods [124, 48, 125, 126], Bayesian networks [127, 128], and deep neural network methods [129, 130]. Considering the taxonomy rules of RSISR methods in this paper, the survey of deep learning-based RSISR methods are missing in this section, which can be found in the corresponding sections later. Specifically, the detailed investigations on these types of RSISR methods are as follows:
@@
Kernel Regression. In 2010, Kim and Kwon [131] representatively proposed a super-resolution method based on kernel ridge regression (KRR), which uses the ${\ell^{2}}$-norm to train and optimize the mapping regression function. KRR is one of the simplest mapping functions for SR tasks. Furthermore, more complex super-resolution schemes have been proposed [132]. In addition, Zhou et al. [125] proposed a novel single image super-resolution method that uses kernel hill regression to connect LR image patches based on HR coding coefficients. A non-local steering kernel regression (NLSKR) method was used by Zhang et al. [133] to perform single image super-resolution tasks, which designed an effective regularization term based on the complementary properties of local structural regularity and non-local self-similarity of images, aiming to preserve sharp edges and produce fine details in the generated images. Similar works can also be found in their previous studies [134]. In RSISR task, a recent work proposed a new synthetic aperture radars (SARs) images SR method by Kanakaraj et al. [135] which extended kernel regression to the adaptive importance sampling unscented Kalman filter (AISUKF) framework to deal with speckle noise. Besides, a robust locally weighted regression method proposed by Ma et al. [47] also uses the kernel regression to perform RSISR task.
@@
Support Vector Regression (SVR). As a kernel regression method, the SVR-based RSISR method has also been widely studied. More concretely, SVR methods involve using high-dimensional feature space to find nonlinear mappings, which is widely used for RSISR tasks and achieves good performance. For example, Zhang et al. [124] proposed a SVR-based RSISR method that learns prior knowledge between HR and LR images. In particular, bilinear interpolation is employed in their method to upsample to the same resolution and preserve high-frequency information. Besides, Zhang et al. [48] proposed an SVR-based super-resolution mapping method (SVR_SRM) for the land cover map. Specifically, SVR_SRM learns the nonlinear relationship between coarse-resolution pixels and corresponding labeled sub-pixels to generate fine-resolution land cover maps from the selected best matching training data. However, SVR_SRM method is sensitive to fraction errors, resulting in many linear artifacts and speckles in the super-resolution results. To this end, Zhang et al. [136] again proposed an improved SVR_SRM method, which mainly used the back-projection operation and the local smoothness prior model to improve the performance of RSISR.
@@
General Regression Neural Network (GRNN). General Regression Neural Networks (GRNN) is a memory-based neural network model [137], which has been used in some studies to improve the performance of RSISR. For example, Li et al. [51] proposed a new GRNN-based method for super-resolution mapping of urban flooding (SMUF) in remote sensing images. In the GRNN-SMUF method [51], a local SMUF model is constructed based on GRNN to describe the relationship between the sub-pixel distribution within a mixed pixel and the scores of the eight neighboring pixels of the mixed pixel. In 2019, Li et al. [59] once again proposed a new SMUF method based on support vector machine and general regression neural network, and named it FSVMGRNN-SMUF. Different from GRNN-SMUF, the FSVMGRNN-SMUF method uses a combination of several SVMs to build a local SMUF model and achieves better performance.
@@
Other Regression Methods. In addition to the above-mentioned regression-based methods, some other regression-based RSISR approaches have also been proposed. For example, Yang et al. [138] trained several simple mapping functions to perform super-resolution tasks. In additoin, Dai et al. [49] proposed an adaptive optimal regressor selection method for the super-resolution task, which selects the most appropriate regressor for each input patch, which jointly produces the minimum super-resolution error for all training data. Yang et al. [139] proposed a soft-assignment based multi-regression method for RSISR, which reconstructs a high-resolution (HR) patch through a dictionary corresponding to its K-nearest training patch clusters. Gao et al. [140] proposed a self-dictionary regression based method for hyperspectral image super-resolution.
@@ 3.2 Deep learning-based RSISR Methods
In this section, supervised deep learning-based RSISR methods are surveyed and discussed in detail from the perspectives of CNN-based, GAN-based, Transformer-based, Mamba-based, and hybrid methods. More specifically, in order to provide a comprehensive and in-depth discussion of the methods in this category, we conducted a detailed discussion of the literatures up to 2025.
@@ 3.2.1 CNN-based RSISR
The prototype of convolutional neural network (CNN) originated from LeNet-5 proposed by LeCun et al. [141], which was the first CNN model to be successfully applied to visual recognition. Since 2012, benefiting from the rapid development of computing resources and hardware technologies such as GPU, CNN had been widely studied and many CNN models had been proposed, such as AlexNet [142], VGG [143] and ResNet [144], etc. These CNN architectures are widely used in computer vision tasks such as image segmentation [145], object detection [146] and image enhancement [147], etc.
In recent years, CNN-based super-resolution methods have been widely studied. Dong et al. [32, 33] first applied CNN to the super-resolution task and achieved good performance. In addition, benefiting from the great success of deep learning, many CNN-based RSISR methods have also been proposed. To the best of our knowledge, Leibel and Korner [148] proposed the first CNN-based RSISR method, namely msiSRCNN, which utilizes remote sensing data (Sentinel-2 images) to train and fine-tune SRCNN [32, 33], as shown in Fig. 6. Specifically, they fine-tuned SRCNN based on the luminance component of the YCbCr color space of remote sensing images and performed performance evaluation.
@@
Multi-Scale Features. Many studies utilized multi-scale features to improve the performance of RSISR. For example, Fu et al. [149] proposed a multi-scale CNN framework (RSCNN) for the RSISR task. Specifically, RSCNN employed convolution kernels of different sizes to extract multi-scale features of images and improved the reconstruction performance of HR images. The idea of extracting multi-scale features based on convolution kernels of different sizes is also involved in these works [150, 151]. A multi-scale residual neural network (MRNN) proposed by Lu et al. [152] also uses complementary features in the patches of different sizes form LR satellite images. Dong et al. [153] proposed a second-order multi-scale super-resolution network (SMSR) for RSISR. In addition, Wang et al. [154] proposed an adaptive multi-scale fusion network (AMFFN) for RSISI. Specifically, several adaptive multi-scale feature extraction (AMFE) modules and the squeeze-and-excited adaptive gating modules are utilized for feature extraction and fusion. In the work of Huan et al. [155], a symmetric multi-scale super-resolution network (AMSSRN) was designed to fully extract and utilize image features. Specifically, they proposed a residual multi-scale block (RMSB) and a residual multi-scale erosion block (RMSDB) to extract shallow and deep features of the image. Multi-scale feature representation is also involved in the RSISR work of Jiang et al. [156], who proposed a deep distillation recurrent network (DDRN) that performs multi-path feature sharing through ultra-dense residual blocks (UDBs) and multi-scale purification unit (MSPU). Besides, DeepNESRM method [157] uses multi-level feature fusion CNN to eliminate score errors while combining multi-scale information to optimize the mixed pixel problem. On the other hand, residual connections play an important role in solving the gradient vanishing problem of deep networks and are widely used in RSISI tasks. For examplel, Huan et al. [158] proposed a new pyramidal multi-scale residual network (PMSRN) for RSISR, which adopts hierarchical residual-like connections to imporve the performance. Deeba et. al [159] proposed a transferred wide residual deep neural network model for RSISR, which reduces the memory cost of the network by increasing the width of the residual network and reducing its depth. In addition, Ye et al. [160] proposed a lightweight RSISR framework, which involves a novel multi-scale residual attention information distillation group to extract richer regional features. More specifically, they adopted a high- and low-frequency separation reconstruction strategy to enhance the high-frequency details of the image. Considering that most CNN methods ignore low-weight background features, Wu et al. [161] proposed a background-based multi-scale feature enhancement network (BMFENet) for RSISR, which constructs a large kernel feature supplement block (LFSB) to generate background feature weights for enhancing the attention of ignored information. Similar work includes the multiscale residual dense network (MRDN) proposed by Kong et al. [162] and the dual-path feature reuse multi-scale network (DFMNet) proposed by Xiao et al. [163].
@@
Similar RSISR methods include Haut et al. [61], which focuses on valuable high-frequency information and ignores uninformative low-frequency features based on the attention mechanism. Considering the necessity of describing remote sensing images from multiple scenes, Zhang et al. [65] proposed a multi-scale attention network (MSAN) for RSISR, which employs a scene-adaptive super-resolution strategy to more accurately describe the structural characteristics of different scenes. In addition, Chen et al. [164] proposed a novel residual split attention network (RSAN) for RSISR, which can simultaneously maintain the overall structure and local details of the image, as shown in Fig. 5 (a). Recently, the research on digital elevation model super-resolution (DEM-SR) has also been carried out, which is of great significance for accurately providing remote sensing geographic information. For example, Chen et al. [165] proposed an integrating attention mechanism (IAM) network method for DEM-SR, which can well capture long-range geographic feature dependencies and local context information, as shown in Fig. 5 (b). In order to solve the problem of limited global receptive field in RSISR, some studies have been performed. For instance, Wang et al. [66] proposed a non-local up-down convolutional attention network (NLASR), which mainly designed a feature enhancement module (NLEB) to model spatial long-range dependencies. Wang et al. [83] proposed a hybrid attention U-type network (HAUNet) for the RSISR task to fuse multi-scale global semantic information. In addition, some research works are devoted to solving the problem of high- and low-frequency information aliasing in CNN-based RSISR methods. For example, Peng et al. [166] designed a gated convolution unit to separate the frequency domain information flow, and further proposed a pre-training of gated convolution neural network (PGCNN) for RSISR. Besides, Zhao et al. [167] used a structure-texture dual preservation strategy to constrain edge generation. Zhang et al. [77] further preposed a residual balanced attention network (RBAN) and combined adversarial training to improve the authenticity of high-frequency textures. Aiming at the large-scale super-resolution reconstruction of remote sensing images, Gao et al. [69] proposed a dense sampling super-resolution network (DSSR) RSISR method, which introduced a dense sampling mechanism, a wide feature attention module and a chain training strategy to improve the network performance. For multi-level feature fusion, Wang et al. [168] proposed a multi-scale spatial-spectral network (M3SN) method, which couples 3D convolution with spectral attention, aiming to deeply mine shallow and deep spatial-spectral features in hyperspectral image super-resolution tasks, thereby improving reconstruction performance. With similar motivations, Chen et al. [169] proposed a residual aggregation-segmentation attention fusion network (RASAF) to achieve cross-channel interaction for RSISR.
@@
On the other hand, some attention-based strategies have been studied to improve the local detail recovery capability of RSISR methods. For example, Ma et al. [170] proposed a dense channel attention mechanism to enhance high-frequency feature reuse. Patnaik et al. [171] combined multi-scale residual attention to capture fine-grained texture for RSISR, and Zhang et al. [172] introduced hybrid high-order attention (HOA) to mine deep statistical associations. In addition, Huang et al. [173] proposed a deep residual dual attention network (DRDAN) method, which introduced a residual dual attention module (RDAB) and a local multi-level fusion module to fuse global and local information. Wang et al. [76] proposed a lightweight feature enhancement network (FeNet) for RSISR, which constructed a lightweight lattice block (LLB) and used channel separation and weight sharing to reduce the computation amount. In the field of lightweight research, Wang et al. [174]proposed a feedback ghost residual dense network (FGRDN), which introduces a spatial-channel attention module (SCM) to extract key features and reduces redundant parameters through feedback mechanism and Ghost module. The enhanced residual convolutional neural network (ERCNN) proposed by Ren et al. [175] also uses a feature attention module, which enhances the discriminative learning ability across feature maps and combines a dual-brightness scheme to optimize high-frequency recovery.
@@
Transform domain. In recent years, some studies have combined CNN with transform domain to perform RSISR, which utilizes the frequency band decomposition capability of domain transform (such as wavelet transform and Fourier transform) and the feature modeling advantages of deep learning to improve the reconstruction performance. For example, Wang et al. [176] proposed a RSISR method combining wavelet multi-scale decomposition with multi-branch CNN. Specifically, they trained the network independently and predicted sub-images in different directions and frequency bands respectively, and then synthesized high-resolution aerial images through inverse wavelet transform, which verified the applicability of frequency domain modeling for complex scenes. Two studies in 2019 further deepened the frequency domain framework in RSISR. Ma et al. [16] proposed a recursive residual network (Recursive Res-Net) to predict high-frequency components, innovatively replaced low-frequency subbands with low-resolution images to preserve details, and optimized network efficiency by removing batch normalization layers. In addition, Yang et al. [177] proposed a multi-scale wavelet 3D convolutional neural network (MW-3D-CNN) for hyperspectral image super-resolution, which uses three-dimensional convolution to extract spatial spectral features and uses L1 loss to optimize multi-subband wavelet coefficient prediction, realizing joint spatial-spectral reconstruction of hyperspectral data. Furthermore, the work of Aburaed et al. [178] also verified the effectiveness of wavelet analysis for reconstructing high-frequency details. In 2021, Deeba et al. [179] proposed a fast wavelet super-resolution framework (FWSR), which accelerates image restoration through approximate sub-band input and sub-pixel layer, and enhances the ability to retain high-frequency features. Sandana et al. [180] proposed a CNN combining wavelet and multi-wavelet transform (CWM-CNN), which integrates discrete wavelet and multi-wavelet transform to extract shallow features, and combines a compact three-layer CNN to achieve parameter-efficient high-resolution reconstruction. Recently, Wang et al. [92] proposed a two-stage joint space-frequency network (TSFNet) to address the problem of large-scale super-resolution. Specifically, they designed an amplitude-guided phase adaptive filtering module (AGPF) to decouple frequency domain degradation features, and achieved progressive detail optimization through cross-stage feature fusion, achieving good performance in maintaining structural integrity.
@@
Spatial-Spectral Joint Modeling. In recent years, some studies have focused on joint spatial-spectral modeling to achieve higher quality reconstruction by fusing the spatial context and spectral correlation of images. For example, early work [181] employed CNN to directly learn the mapping from low-resolution bands to high-resolution bands end-to-end based on the inherent resolution differences of Sentinel-2 multi-bands (10m/20m/60m), while implicitly utilizing the spatial features of high-resolution bands to guide the reconstruction of low-resolution bands. Yin et al. [182] proposed a DeepRivSRM method for RSISR. Specifically, they extracted the river spectral features in mixed pixels by spectral unmixing, and then combined the spatial context for super-resolution mapping. To further enhance physical consistency, some studies have injected explicit physical model constraints. For example, Arun et al. [183] proposed a collaborative unmixing framework based on sparse coding for RSISR, which jointly optimized the sparse dictionary and spectral reconstruction, and fused the spatial-spectral prior through the encoding-decoding architecture to ensure spectral fidelity. In addition, Wagner et al. [184] employed a deep residual network (ResNet) to fuse shallow spectral features with deep spatial features via jump connections to achieve end-to-end spectral-spatial joint optimization. Vasilescu et al. [185] innovatively embedded the modulation transfer function (MTF) of the sensor into the multi-objective loss function to simulate the sensor degradation process and force the network to learn the physical consistency of the inverse mapping. In addition, they combined high-frequency detail alignment loss (such as 10m band gradient matching) to improve the accuracy of spatial detail reconstruction. In the recent work of Chen et al. [186], a spatial-spectral attention module (DSSA) combined with a dilated convolution method was proposed to extract global spatial features, and the 3D Inception module was used to mine multi-scale spectral information to solve the problem of spectral-spatial information imbalance in hyperspectral images. In addition, Mei et al. [187] proposed two end-to-end joint optimization framework for RSISR, namely simultaneous spatial–spectral joint SR (SimSSJSR) and separated spatial–spectral joint SR (SepSSJSR). Specifically, spectral unmixing, regularized deconvolution, and endmember similarity constraints are employed to ensure spectral consistency. Müller et al. [188] performed panchromatic band (PAN) and multispectral band fusion, which used CNN training to improve the spatial resolution of multispectral satellite images.
@@
Local-Global Feature Fusion. This type of methods usually captures the local detail features and global prior information of remote sensing images, and uses multi-scale feature fusion mechanism to improve the quality of super-resolution reconstruction. For example, the local–global combined networks (LGCNet) proposed by Lei et al. [53] adopts a multi-branch structure to achieve multi-level representation learning. Besides, Xu et al. [189] proposed a deep memory connected network (DMCN) for RSISR, which achieves feature integration by building local-global memory connections. These studies all emphasize the integration of scene-level semantic information while maintaining local texture details. They generally introduce spatial downsampling units to optimize computational efficiency and balance model complexity and performance by compressing the size of feature maps.

### v3

@@ 3.1.3 Regression-based Methods
The regression-based RSISR methods learn the function mappings between LR and HR images, which regard super-resolution as a regression problem. More specifically, these methods usually extract features from LR images based on paired datasets to train regression models, and then use the trained models to predict and reconstruct HR images. Typically, regression-based methods consist of three main steps: regression model training, projecting of LR images to HR image space, and final super-resolution result generation, as shown in Fig. 6.
Figure 6: The schematic diagram of the regression-based RSISR method.
@@
In the past two decades, various regression-based super-resolution methods have been proposed, including kernel regression-based methods [126, 49, 127, 128], Bayesian networks [129, 130], and deep neural network methods [131, 132]. Considering the taxonomy rules of RSISR methods in this paper, the survey of deep learning-based RSISR methods is missing in this section, which can be found in the corresponding sections later. Specifically, the detailed investigations on these types of RSISR methods are as follows:
@@
Kernel Regression. In 2010, Kim and Kwon [133] representatively proposed a super-resolution method based on kernel ridge regression (KRR), which uses the $\ell^{2}$}-norm to train and optimize the mapping regression function. KRR is one of the simplest mapping functions for SR tasks. Furthermore, more complex super-resolution schemes have been proposed [134]. In addition, Zhou et al. [127] proposed a novel single-image super-resolution method that uses kernel hill regression to connect LR image patches based on HR coding coefficients. A non-local steering kernel regression (NLSKR) method was used by Zhang et al. [135] to perform single image super-resolution tasks, which designed an effective regularization term based on the complementary properties of local structural regularity and non-local self-similarity of images, aiming to preserve sharp edges and produce fine details in the generated images. Similar works can also be found in their previous studies [136]. In the RSISR task, a recent work proposed a new synthetic aperture radars (SARs) image SR method by Kanakaraj et al. [137] which extended kernel regression to the adaptive importance sampling unscented Kalman filter (AISUKF) framework to deal with speckle noise. Besides, a robust locally weighted regression method proposed by Ma et al. [48] also uses the kernel regression to perform RSISR task.
@@
Support Vector Regression (SVR). As a kernel regression method, the SVR-based RSISR method has also been widely studied. More concretely, SVR methods involve using a high-dimensional feature space to find nonlinear mappings, which is widely used for RSISR tasks and achieves good performance. For example, Zhang et al. [126] proposed an SVR-based RSISR method that learns prior knowledge between HR and LR images. In particular, bilinear interpolation is employed in their method to upsample to the same resolution and preserve high-frequency information. Besides, Zhang et al. [49] proposed an SVR-based super-resolution mapping method (SVR_SRM) for the land cover map. Specifically, SVR_SRM learns the nonlinear relationship between coarse-resolution pixels and corresponding labeled sub-pixels to generate fine-resolution land cover maps from the selected best-matching training data. However, SVR_SRM method is sensitive to fraction errors, resulting in many linear artifacts and speckles in the super-resolution results. To this end, Zhang et al. [138] again proposed an improved SVR_SRM method, which mainly used the back-projection operation and the local smoothness prior model to improve the performance of RSISR.
@@
General Regression Neural Network (GRNN). General Regression Neural Networks (GRNN) is a memory-based neural network model [139], which has been used in some studies to improve the performance of RSISR. For example, Li et al. [52] proposed a new GRNN-based method for super-resolution mapping of urban flooding (SMUF) in RSIs. In the GRNN-SMUF method [52], a local SMUF model is constructed based on GRNN to describe the relationship between the sub-pixel distribution within a mixed pixel and the scores of the eight neighboring pixels of the mixed pixel. In 2019, Li et al. [60] once again proposed a new SMUF method based on a support vector machine and a general regression neural network, and named it FSVMGRNN-SMUF. Different from GRNN-SMUF, the FSVMGRNN-SMUF method uses a combination of several SVMs to build a local SMUF model and achieves better performance.
@@
Other Regression Methods. In addition to the above-mentioned regression-based methods, some other regression-based RSISR approaches have also been proposed. For example, Yang et al. [140] trained several simple mapping functions to perform super-resolution tasks. In addition, Dai et al. [50] proposed an adaptive optimal regressor selection method for the super-resolution task, which selects the most appropriate regressor for each input patch, jointly producing the minimum super-resolution error for all training data. Yang et al. [141] proposed a soft-assignment based multi-regression method for RSISR, which reconstructs a high-resolution (HR) patch through a dictionary corresponding to its K-nearest training patch clusters. Gao et al. [142] proposed a self-dictionary regression-based method for hyperspectral image super-resolution.
@@ 3.2 Deep learning-based RSISR Methods
In this section, supervised deep learning-based RSISR methods are surveyed and discussed in detail from the perspectives of CNN-based, GAN-based, Transformer-based, Mamba-based, and hybrid methods. More specifically, in order to provide a comprehensive and in-depth discussion of the methods in this category, we conducted a detailed discussion of the literature up to 2025.
@@ 3.2.1 CNN-based RSISR
The prototype of a convolutional neural network (CNN) originated from LeNet-5 proposed by LeCun et al. [143], which was the first CNN model to be successfully applied to visual recognition. Since 2012, benefiting from the rapid development of computing resources and hardware technologies such as GPU, CNN has been widely studied and many CNN models have been proposed, such as AlexNet [144], VGG [145], ResNet [146], and OverLoCK [147], etc. These CNN architectures are widely used in computer vision tasks such as image segmentation [148], object detection [149], and image enhancement [150], etc.
In recent years, CNN-based super-resolution methods have been widely studied. Dong et al. [32, 33] first applied CNN to the super-resolution task and achieved good performance. In addition, benefiting from the great success of deep learning, many CNN-based RSISR methods have also been proposed. To the best of our knowledge, Leibel and Korner [151] proposed the first CNN-based RSISR method, namely msiSRCNN, which utilizes remote sensing data (Sentinel-2 images) to train and fine-tune SRCNN [32, 33], as shown in Fig. 7. Specifically, they fine-tuned SRCNN based on the luminance component of the YCbCr color space of RSIs and performed a performance evaluation.
@@
Multi-Scale Features. Many studies utilized multi-scale features to improve the performance of RSISR. For example, Fu et al. [152] proposed a multi-scale CNN framework (RSCNN) for the RSISR task. Specifically, RSCNN employed convolution kernels of different sizes to extract multi-scale features of images and improved the reconstruction performance of HR images. The idea of extracting multi-scale features based on convolution kernels of different sizes is also involved in these works [153, 154]. A multi-scale residual neural network (MRNN) proposed by Lu et al. [155] also uses complementary features in the patches of different sizes from LR satellite images. Dong et al. [156] proposed a second-order multi-scale super-resolution network (SMSR) for RSISR. In addition, Wang et al. [157] proposed an adaptive multi-scale fusion network (AMFFN) for RSISI. Specifically, several adaptive multi-scale feature extraction (AMFE) modules and the squeeze-and-excition adaptive gating modules are utilized for feature extraction and fusion. In the work of Huan et al. [158], a symmetric multi-scale super-resolution network (AMSSRN) was designed to fully extract and utilize image features. Specifically, they proposed a residual multi-scale block (RMSB) and a residual multi-scale erosion block (RMSDB) to extract shallow and deep features of the image. Multi-scale feature representation is also involved in the RSISR work of Jiang et al. [159], who proposed a deep distillation recurrent network (DDRN) that performs multi-path feature sharing through ultra-dense residual blocks (UDBs) and a multi-scale purification unit (MSPU). Besides, DeepNESRM method [160] uses a multi-level feature fusion CNN to eliminate score errors while combining multi-scale information to optimize the mixed pixel problem. On the other hand, residual connections play an important role in solving the gradient vanishing problem of deep networks and are widely used in RSISI tasks. For example, Huan et al. [161] proposed a new pyramidal multi-scale residual network (PMSRN) for RSISR, which adopts hierarchical residual-like connections to improve the performance. Deeba et. al [162] proposed a transferred wide residual deep neural network model for RSISR, which reduces the memory cost of the network by increasing the width of the residual network and reducing its depth. In addition, Ye et al. [163] proposed a lightweight RSISR framework, which involves a novel multi-scale residual attention information distillation group to extract richer regional features. More specifically, they adopted a high- and low-frequency separation reconstruction strategy to enhance the high-frequency details of the image. Considering that most CNN methods ignore low-weight background features, Wu et al. [164] proposed a background-based multi-scale feature enhancement network (BMFENet) for RSISR, which constructs a large kernel feature supplement block (LFSB) to generate background feature weights for enhancing the attention of ignored information. Similar work includes the multiscale residual dense network (MRDN) proposed by Kong et al. [165] and the dual-path feature reuse multi-scale network (DFMNet) proposed by Xiao et al. [166].
@@
Similar RSISR methods include Haut et al. [62], which focuses on valuable high-frequency information and ignores uninformative low-frequency features based on the attention mechanism. Considering the necessity of describing RSIs from multiple scenes, Zhang et al. [66] proposed a multi-scale attention network (MSAN) for RSISR, which employs a scene-adaptive super-resolution strategy to more accurately describe the structural characteristics of different scenes. In addition, Chen et al. [167] proposed a novel residual split attention network (RSAN) for RSISR, which can simultaneously maintain the overall structure and local details of the image, as shown in Fig. 4 (a). Recently, research on digital elevation model super-resolution (DEM-SR) has also been carried out, which is of great significance for accurately providing remote sensing geographic information. For example, Chen et al. [168] proposed an integrating attention mechanism (IAM) network method for DEM-SR, which can well capture long-range geographic feature dependencies and local context information, as shown in Fig. 4 (b). In order to solve the problem of the limited global receptive field in RSISR, some studies have been performed. For instance, Wang et al. [67] proposed a non-local up-down convolutional attention network (NLASR), which mainly designed a feature enhancement module (NLEB) to model spatial long-range dependencies. Wang et al. [84] proposed a hybrid attention U-type network (HAUNet) for the RSISR task to fuse multi-scale global semantic information. In addition, some research works are devoted to solving the problem of high- and low-frequency information aliasing in CNN-based RSISR methods. For example, Peng et al. [169] designed a gated convolution unit to separate the frequency domain information flow, and further proposed a pre-training of a gated convolution neural network (PGCNN) for RSISR. Besides, Zhao et al. [170] used a structure-texture dual preservation strategy to constrain edge generation. Zhang et al. [78] further proposed a residual balanced attention network (RBAN) and combined adversarial training to improve the authenticity of high-frequency textures. Aiming at the large-scale super-resolution reconstruction of RSIs, Gao et al. [70] proposed a dense sampling super-resolution network (DSSR) RSISR method, which introduced a dense sampling mechanism, a wide feature attention module, and a chain training strategy to improve the network performance. For multi-level feature fusion, Wang et al. [171] proposed a multi-scale spatial-spectral network (M3SN) method, which couples 3D convolution with spectral attention, aiming to deeply mine shallow and deep spatial-spectral features in hyperspectral image super-resolution tasks, thereby improving reconstruction performance. With similar motivations, Chen et al. [172] proposed a residual aggregation-segmentation attention fusion network (RASAF) to achieve cross-channel interaction for RSISR.
@@
On the other hand, some attention-based strategies have been studied to improve the local detail recovery capability of RSISR methods. For example, Ma et al. [173] proposed a dense channel attention mechanism to enhance high-frequency feature reuse. Patnaik et al. [174] combined multi-scale residual attention to capture fine-grained texture for RSISR, and Zhang et al. [175] introduced hybrid high-order attention (HOA) to mine deep statistical associations. In addition, Huang et al. [176] proposed a deep residual dual attention network (DRDAN) method, which introduced a residual dual attention module (RDAB) and a local multi-level fusion module to fuse global and local information. Wang et al. [77] proposed a lightweight feature enhancement network (FeNet) for RSISR, which constructed a lightweight lattice block (LLB) and used channel separation and weight sharing to reduce the computation amount. In the field of lightweight research, Wang et al. [177]proposed a feedback ghost residual dense network (FGRDN), which introduces a spatial-channel attention module (SCM) to extract key features and reduces redundant parameters through a feedback mechanism and a Ghost module. The enhanced residual convolutional neural network (ERCNN) proposed by Ren et al. [178] also uses a feature attention module, which enhances the discriminative learning ability across feature maps and combines a dual-brightness scheme to optimize high-frequency recovery.
@@
Transform domain. In recent years, some studies have combined CNN with a transform domain to perform RSISR, which utilizes the frequency band decomposition capability of domain transform (such as wavelet transform and Fourier transform) and the feature modeling advantages of deep learning to improve the reconstruction performance. For example, Wang et al. [179] proposed an RSISR method combining wavelet multi-scale decomposition with a multi-branch CNN. Specifically, they trained the network independently and predicted sub-images in different directions and frequency bands, respectively, and then synthesized high-resolution aerial images through inverse wavelet transform, which verified the applicability of frequency domain modeling for complex scenes. Two studies in 2019 further deepened the frequency domain framework in RSISR. Ma et al. [16] proposed a recursive residual network (Recursive Res-Net) to predict high-frequency components, innovatively replaced low-frequency subbands with low-resolution images to preserve details, and optimized network efficiency by removing batch normalization layers. In addition, Yang et al. [180] proposed a multi-scale wavelet 3D convolutional neural network (MW-3D-CNN) for hyperspectral image super-resolution, which uses three-dimensional convolution to extract spatial spectral features and uses L1 loss to optimize multi-subband wavelet coefficient prediction, realizing joint spatial-spectral reconstruction of hyperspectral data. Furthermore, the work of Aburaed et al. [181] also verified the effectiveness of wavelet analysis for reconstructing high-frequency details. In 2021, Deeba et al. [182] proposed a fast wavelet super-resolution framework (FWSR), which accelerates image restoration through approximate sub-band input and sub-pixel layer, and enhances the ability to retain high-frequency features. Sandana et al. [183] proposed a CNN combining wavelet and multi-wavelet transform (CWM-CNN), which integrates discrete wavelet and multi-wavelet transform to extract shallow features, and combines a compact three-layer CNN to achieve parameter-efficient high-resolution reconstruction. Recently, Wang et al. [93] proposed a two-stage joint space-frequency network (TSFNet) to address the problem of large-scale super-resolution. Specifically, they designed an amplitude-guided phase adaptive filtering module (AGPF) to decouple frequency domain degradation features and achieved progressive detail optimization through cross-stage feature fusion, achieving good performance in maintaining structural integrity.
@@
Spatial-Spectral Joint Modeling. In recent years, some studies have focused on joint spatial-spectral modeling to achieve higher quality reconstruction by fusing the spatial context and spectral correlation of images. For example, early work [184] employed CNN to directly learn the mapping from low-resolution bands to high-resolution bands end-to-end based on the inherent resolution differences of Sentinel-2 multi-bands (10m/20m/60m), while implicitly utilizing the spatial features of high-resolution bands to guide the reconstruction of low-resolution bands. Yin et al. [185] proposed a DeepRivSRM method for RSISR. Specifically, they extracted the river spectral features in mixed pixels by spectral unmixing and then combined the spatial context for super-resolution mapping. To further enhance physical consistency, some studies have injected explicit physical model constraints. For example, Arun et al. [186] proposed a collaborative unmixing framework based on sparse coding for RSISR, which jointly optimized the sparse dictionary and spectral reconstruction, and fused the spatial-spectral prior through the encoding-decoding architecture to ensure spectral fidelity. In addition, Wagner et al. [187] employed a deep residual network (ResNet) to fuse shallow spectral features with deep spatial features via jump connections to achieve end-to-end spectral-spatial joint optimization. Vasilescu et al. [188] innovatively embedded the modulation transfer function (MTF) of the sensor into the multi-objective loss function to simulate the sensor degradation process and force the network to learn the physical consistency of the inverse mapping. In addition, they combined high-frequency detail alignment loss (such as 10m band gradient matching) to improve the accuracy of spatial detail reconstruction. In the recent work of Chen et al. [189], a spatial-spectral attention module (DSSA) combined with a dilated convolution method was proposed to extract global spatial features, and the 3D Inception module was used to mine multi-scale spectral information to solve the problem of spectral-spatial information imbalance in hyperspectral images. In addition, Mei et al. [190] proposed two end-to-end joint optimization frameworks for RSISR, namely simultaneous spatial–spectral joint SR (SimSSJSR) and separated spatial–spectral joint SR (SepSSJSR). Specifically, spectral unmixing, regularized deconvolution, and endmember similarity constraints are employed to ensure spectral consistency. Müller et al. [191] performed panchromatic band (PAN) and multispectral band fusion, which used CNN training to improve the spatial resolution of multispectral satellite images.
@@
Local-Global Feature Fusion. This type of method usually captures the local detail features and global prior information of RSIs, and uses a multi-scale feature fusion mechanism to improve the quality of super-resolution reconstruction. For example, the local–global combined networks (LGCNet) proposed by Lei et al. [54] adopts a multi-branch structure to achieve multi-level representation learning. Besides, Xu et al. [192] proposed a deep memory connected network (DMCN) for RSISR, which achieves feature integration by building local-global memory connections. These studies all emphasize the integration of scene-level semantic information while maintaining local texture details. They generally introduce spatial downsampling units to optimize computational efficiency and balance model complexity and performance by compressing the size of feature maps.